Circular No. 6159

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NEARBY GALAXIES MB 1 AND MB 2 M. L. McCall, York University; R. J. Buta, University of Alabama; and

W. K. Huchtmeier, Max-Planck-Institut für Radioastronomie; write: “We report on the discovery of two new nearby galaxies likely to be associated with the heavily-obscured giant elliptical galaxy Maffei 1. The two galaxies were discovered (by McCall and Buta) on deep I-band CCD images of a 1^{o} field centered near Maffei 1 and acquired with the Kitt Peak Burrell Schmidt telescope on 1992 Oct. 25. The larger of the two galaxies (MB 1) is located at α = 2^{h}31^{m}52^{s}.1, δ = +59^{o}09′39′′.7 (equinox 1950.0), which is only 18′ to the southwest of Maffei 1. It is a late-type spiral (probably Sd) with a strong oval distortion, possibly as big as 6′ across; due to the low apparent surface brightness, the inclination is uncertain. The smaller of the two new galaxies (MB 2) is located at α = 2^{h}33^{m}16^{s}.1, δ = +59^{o}01′12′′.8, which is 25′ nearly due south of Maffei 1. It is probably a dwarf irregular galaxy, the diameter being 1′.4. The optical observations will appear in the June 1995 issue of the A.J. MB 1 was detected (by Huchtmeier) at 21 cm with the 100m Effelsberg telescope on 1995 Mar. 2; the line is centered on a heliocentric velocity of 189 ± 3 km/s, and displays a peak flux of 0.16 ± 0.01 Jy and an integrated flux of 9.0 Jy km/s. The profile is nearly gaussian, with full widths of 85, 79, and 60 km/s at 20, 25, and 50 percent, respectively, of the peak. The radial velocity and proximity to Maffei 1 on the sky indicate that this object is a member of a group that is 2 to 4 Mpc distant and is now known to be composed of 13 galaxies that are dominated by Maffei 1 and IC 342. For its area, MB 1 is deficient in hydrogen, suggesting that it has directly interacted with Maffei 1. MB 2 was not detected at 21 cm over a search band from −500 to +4000 km/s. Weak extragalactic emission could be hidden within the range of local hydrogen (i.e., −140 to +10 km/s) or in an interference zone around 1500 to 1600 km/s. Again, the proximity to Maffei 1 suggests an association with that galaxy.”

AL COMAE BERENICES Further visual magnitude estimates (cf. IAUC 6155): Apr. 7.22 UT,

12.3 (W. G. Dillon, Missouri City, TX); 8.032, 12.7 (T. Vanmunster, Landen, Belgium); 8.912, 12.6 (B. H. Granslo, Fjellhamar, Norway); 9.826, 12.7 (P. Schmeer, Bischmisheim, Germany).

1995 April 11 Daniel W. E. Green